OpenStudy (anonymous):

x2 + y2 = 8 x - y = 0 Select all of the following that are solutions to the system shown. (2, 2) (2, -2) (-2, 2) (-2, -2)

OpenStudy (anonymous):

http://openstudy.com/study#/updates/51d338b9e4b0befebdaf0cfe I need help please reply to this one also.

OpenStudy (anonymous):

To determine if the point is in the system, we simply plug in to the equations and see if they work. Lets work with the second one to narrow it down. Add y to both sides to get \[x=y\] We can eliminate the second and third options now because the x isnt' equal to the y. Now plug those two into the first equation to check them.

OpenStudy (anonymous):

x - y = 0 x=y a and d then

OpenStudy (anonymous):

It's the first option (2,2) Because when 2^2+2^2=8 4+4=8 8=8 And when x - y = 0 2-2=0 0=0 I proved my answer xD